### v1

# Soteria: Language-Specific Functional Parameter Steering for Multilingual Safety Alignment

Abstract
Ensuring consistent safety across multiple languages remains a significant challenge for large language models (LLMs). We introduce Soteria, a lightweight yet powerful strategy that locates and minimally adjusts the “functional heads” most responsible for harmful content generation in each language. By altering only a fraction of parameters, Soteria drastically reduces policy violations without sacrificing overall model performance, even in low-resource settings. To rigorously evaluate our approach, we also present XThreatBench, a specialized multilingual dataset capturing fine-grained harmful behaviors drawn from real policy guidelines. Experiments with leading open-source LLMs (e.g., Llama, Qwen, Mistral) show that Soteria consistently improves safety metrics across high-, mid-, and low-resource languages. These findings highlight a promising path toward scalable, linguistically attuned, and ethically aligned LLMs worldwide. The source code and dataset are available at: https://github.com/NeuralSentinel/Soteria
[…]
[ topline=false, bottomline=false, skipabove=skipbelow=leftline=true, rightline=true, linecolor=NavyBlue, linewidth=2pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!10, roundcorner=10pt ]stylishframe
Soteria: Language-Specific Functional Parameter Steering for Multilingual Safety Alignment
Somnath Banerjee, Sayan Layek, Pratyush Chatterjee, Animesh Mukherjee, Rima Hazra Indian Institute of Technology Kharagpur INSAIT, Sofia University “St. Kliment Ohridski”
1 Introduction
LLMs such as GPT-4o OpenAI et al. (2024), Claude Anthropic , and Llama Touvron et al. (2023) have revolutionized the AI landscape by delivering impressive performance across tasks ranging from text generation to question answering. These breakthroughs stem from extensive pre-training on large, diverse corpora Zhou et al. (2023); Kamalloo et al. (2023); Nguyen et al. (2024a). Yet, much of the early research on LLMs’ multilingual capabilities relied on translating English queries into non-English, a strategy that obscures genuine multilingual performance Zhao et al. (2024). Although newer LLMs feature advanced tokenizers that handle non-English inputs more effectively, key safety measures, including red teaming and content filtering remain predominantly English centric Zhang et al. (2023); Gurgurov et al. (2024).
As a result, non-English use cases are comparatively under-protected, and especially smaller-parameter models (e.g., 8B or 7B) often implemented in low-resource settings are at greater risk of generating harmful or culturally insensitive outputs Banerjee et al. (2025). Moreover, prior work on safety mechanisms has focused mainly on English, overlooking the nuances and needs of broader linguistic communities Banerjee et al. (2024b); Hazra et al. (2024a). In this context, it becomes clear that robust, multilingual safety protocols are essential to protect users and maintain linguistic sensitivity across the globe Wang et al. (2024); Lu and Koehn (2024).
A major obstacle to robust multilingual safety lies in the limitations of early tokenizers Petrov et al. (2023); Hong et al. (2024), which were not designed properly to capture the rich morphological and script diversity in global languages Ali et al. (2024). As a result, LLMs built on these tokenizers struggle to generate linguistically relevant and accurate outputs in non-English settings, undermining the effectiveness of any safety measures. While newer models incorporate more sophisticated multilingual tokenizers [^1] https://huggingface.co/blog/llama31, prior efforts largely treated multilingual support as an afterthought added later via fine-tuning rather than integrated as a core capability Richburg and Carpuat (2024). This approach often relies on “bridging strategies,” such as translating queries into English before applying moderation filters, a practice that can distort content classification Bang et al. (2023); Lai et al. (2024). Even extensive fine-tuning typically fails to address deeper, English-dominant architectural constraints, especially for languages with multiple scripts or highly complex morphology. Moreover, creating large-scale multilingual datasets for each fine-tuning cycle is prohibitively expensive and time-intensive Yu et al. (2022). Although scaling up to larger-parameter models can bolster multilingual proficiency, such approaches may be infeasible in low-resource or time-sensitive contexts Nguyen et al. (2024b); Chelombitko et al. (2024).
[…]
  • To the best of our knowledge, we are the first to introduce a multilingual parameter-efficient safety mechanism – Soteria – that modifies only about $\sim$3of the model’s language-specific “functional heads,” effectively reducing harmful outputs without compromising overall performance.
  • We introduce XThreatBench, a multilingual dataset covering harm categories derived from Meta’s content guidelines, closing critical gaps in existing safety benchmarks.
  • Our experiments encompass a broad linguistic spectrum from high- to low-resource to demonstrate that these safety enhancements are not confined to English or high-resource settings.
2 Related work
Mechanistic interpretability: This section explores how internal LLM components (neurons, layers, attention heads) shape model behaviors Geiger et al. (2021); Stolfo et al. (2023); Gurnee et al. (2023). Early work identified key neurons Hendrycks (2023); Chen et al. (2024), but recent studies underscore attention heads’ critical roles in various language tasks Vig (2019); Wu et al. (2025). Ablation approaches reveal certain heads are crucial for syntactic parsing and factual reasoning Michel et al. (2019); Meng et al. (2023), yet their safety implications remain underexplored Gould et al. (2023); Wang et al. (2023). This gap highlights the need for fine-grained analysis to enhance transparency and safety.
Safety alignment: Efforts to ensure LLM safety focus on mitigating adversarial prompts Xie et al. (2018), designing robust filtering Xiao et al. (2024), and maintaining dynamic oversight Kenton et al. (2024); Wang et al. (2024). Early studies Yao et al. (2024) expose key vulnerabilities and propose ethical risk frameworks. Subsequent work Sachdeva et al. (2025); Banerjee et al. (2024a) reveals how subtle prompt manipulations can evade safeguards, prompting research into attack strategies Wolf et al. (2024) and defenses like RAIN Li et al. (2023). Others emphasize dynamic monitoring Bhardwaj et al. (2024) and adaptive safety mechanisms, including safety arithmetic Hazra et al. (2024b) for test-time alignment and SafeInfer Banerjee et al. (2024b), SafeDecoding Xu et al. (2024) for decoding-time alignment.
3 Methodology
In this section, we present our methodology for identifying and mitigating harmful behavior in LLMs. We first introduce the underlying components of autoregressive LLMs (Section 3.1), focusing on their transformer decoder layers and attention mechanisms. We then describe our framework (Section 3.2) for identifying important attention heads that are crucial for task-solving and language-specific processing, followed by the procedure to remove harm-inducing directions from these heads.
3.1 Preliminaries
We define an autoregressive LLM as $\mathcal{M}$, which comprises multiple transformer decoder layers, denoted by $\mathcal{L}$. Each transformer decoder layer consists of two fundamental modules – multi-head attention ($MHA$) and feed-forward network ($FFN$). The outputs of $MHA$ and $FFN$ modules in layer $l\in\mathcal{L}$ are denoted by $atn^{l}$ and $mlp^{l}$, respectively. The hidden state of a transformer decoder layer $l$ is denoted by $ht_{l}$. The hidden state $ht_{l}$ is computed as shown in Equation 1 where $ht_{l-1}$ represents the hidden state from the previous layer $l-1$.
Figure 1: Schematic diagram of the Soteria.
$\vspace{-0.2cm}ht_{l}=ht_{l-1}+mlp^{l}+atn^{l}$ (1)
Mathematically, the output $atn^{l}$ of $MHA$ module is further obtained using Equation 2 in which each attention head is represented as $h_{i}^{l}$ where $i\in\mathcal{I}$ denotes the $i^{th}$ attention head and $|\mathcal{I}|$ denotes the number of heads in each layer $l$. $W^{O}_{l}\in\mathbb{R}^{|\mathcal{I}|\cdot d_{k}\times d_{m}}$ projects ($O$ Projection) the concatenated heads to the model dimension whereby the head $h_{i}^{l}$ has a dimension of $d_{k}$ and the hidden dimension of the model is $d_{m}$. Each head $h_{i}^{l}$ is derived as given in Equation 3 in which $W_{i}^{Q}$, $W_{i}^{K}$ and $W_{i}^{V}$ denote the learned weight matrices for the query $Q$, key $K$, and values $V$ of the $i^{th}$ head.
$\vspace{-0.2cm}atn_{l}=\text{concat}(h_{1}^{l},\dots,h_{\mathcal{I}}^{l})\cdot W^{O}_{l}$ (2)
$h_{i}^{l}=\text{attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (3)
In this work, similar to Todd et al. (2024), we adopt the attention definition proposed by Elhage et al. (2021) rather than the one introduced in Vaswani et al. (2017). The study in  Elhage et al. (2021) highlights that the formulation in  Vaswani et al. (2017) can be interpreted as decomposing weight matrix $W^{O}_{l}$ into a block form $[W_{l1}^{O}\;W_{l2}^{O}\;\dots\;W_{l\mathcal{I}}^{O}]$, allowing $h_{i}^{l}$ to be directly projected into residual stream space. Each block $W_{li}^{O}\in\mathbb{R}^{d_{k}\times d_{m}}$ determines how information from $h_{i}^{l}$ is transformed into the final model dimension. We use the output $atn_{i}^{l}$ corresponding to $i^{th}$ head as written in Equation 4.
$\vspace{-0.2cm}atn_{i}^{l}=h_{i}^{l}\cdot W_{li}^{O}\in\mathbb{R}^{d_{m}}$ (4)
In this study, we consider a set of languages $\ell\in\mathscr{L}$. To identify important attention heads for each language $\ell$, we define a set of tasks, denoted by $t\in\mathcal{T}$, specific to each language. To mitigate harmful direction, we fine-tune a language model with the same backbone as $\mathcal{M}$ using a dataset $\mathcal{D}_{H}$ consisting of harmful instances resulting in a harmful model $\mathcal{M}_{H}$. The dataset $\mathcal{D}_{H}$ consists of a collection of harmful questions paired with their corresponding harmful answers.
3.2 Our framework
In our framework, we first identify important attention heads (i.e., $atn_{i}^{l}$ for the $i^{th}$ head) and subsequently remove the harm direction from the target model.
Identifying important attention heads: Our objective is to identify attention heads that contribute to both task-solving and language-specific processing. To analyze the role of attention heads in task completion across languages, we translate all tasks into a specific language $\ell$. Unlike prior approaches Tang et al. (2024), we emphasize task relevance to ensure that the identified heads capture task-specific linguistic information. Following Todd et al. (2024), each task $t$ comprises a dataset containing a set of prompts, denoted by $\mathscr{P}^{t}$. A prompt $p_{k}^{t}\in\mathscr{P}^{t}$ is represented as $p_{k}^{t}=\left[(q_{k_{1}},r_{k_{1}}),\cdots,(q_{k_{K}},r_{k_{K}}),q_{k_{Q}}\right]$, where the target answer $r_{k_{Q}}$ for question $q_{k_{Q}}$ is not included in the prompt. Using this prompt $p_{k}^{t}$, the next-token prediction function $\mathcal{M}(p_{k}^{t})$ ranks the correct answer highest, allowing us to assess the contribution of specific attention heads to both task performance and language processing.
We provide the prompt $p_{k}^{t}$ to language model $\mathcal{L}$ so that it can predict the correct answer for the question $q_{k_{Q}}$. Our objective is to identify model components with a causal role in multilingual processing during the prediction of $r_{k_{Q}}$. For each attention head $atn_{i}^{l}$ and task dataset $\mathscr{P^{t}}$, we compute mean condition activations $\hat{atn_{i}^{l}}_{t}$ in Equation 5. In Equation 5, $atn_{i}^{l}(p_{k}^{t})$ is the attention output of prompt $p_{k}^{t}$ for $i^{th}$ attention head.
Figure 2: Identified top 20 heads for Llama 3.1 for Spanish and Bengali.
$\vspace{-0.2cm}\hat{atn_{i}^{l}}_{t}=\frac{1}{|\mathscr{P}_{t}|}\sum_{p_{k}^{t}\in\mathscr{P}_{t}}atn_{i}^{l}(p_{k}^{t})$ (5)
In parallel, we have a corrupted prompt $\hat{p}_{i}^{k}$ where the responses are shuffled $\hat{p}_{i}^{k}=\left[(q_{k_{1}},\hat{r}_{k_{1}}),\cdots,(q_{k_{K}},\hat{r}_{k_{K}}),q_{k_{Q}}\right]$. Next, we pass the corrupted prompt $\hat{p}_{k}^{t}$ through the language model $\mathcal{L}$ and replace a specific attention head activation $atn_{i}^{l}(\hat{p}_{k}^{t})$ with the actual mean task conditioned activation $\hat{atn_{i}^{l}}_{t}$. We attempt to understand how much the actual task conditioned activation can help to predict the correct answer. Further we measure the causal indirect effect (CIE) toward recovering the correct answer $r_{k_{Q}}$ as shown in Equation 6.
$\text{CIE}(atn_{i}^{l}\mid\hat{p}_{k}^{t})$ $=\mathcal{M}\left(\hat{p}_{k}^{t}\mid atn_{i}^{l}:=\hat{atn}_{it}^{l}\right)[r_{k_{Q}}]$ (6)
$\quad-\mathcal{M}(\hat{p}_{k}^{t})[r_{k_{Q}}]$
Further, we obtain the average indirect effect AIE of an attention $atn_{i}^{l}$ ($AIE(atn_{i}^{l})$) by averaging the causal indirect effect across all the tasks and their corrupted prompts. To identify the set of attention heads with the strongest causal effects, we iterate the same process for all the attention heads in the language model $\mathcal{L}$ (see Figure 2). We also repeat the whole process for every language $\ell\in\mathscr{L}$.
Removal of harm direction: According to Equation 4, each block $W_{li}^{O}$ determines the transformation of information from $h_{i}^{l}$ to the output $atn^{l}_{i}$. Given an important attention $atn^{l}_{i}$, we consider the associated block $W_{li}^{O}$ for harm direction removal. We focus solely on the $O$-projection weight, avoiding unnecessary changes to other layer weights, which could compromise the model’s broader capabilities. Following Hazra et al. (2024a) we compute the harm vector $H_{v}$ by taking the element-wise difference between the $\mathcal{M}_{H}$ and $\mathcal{M}$. Further, we keep only those parameters of $H_{v}$ as per selected blocks ($W_{li}^{O}$ for $i^{th}$ head) of the $W^{O}_{l}$ and make the other parameters zero. The harm vector with retained parameters is denoted by $\hat{H}_{v}$. The safe model $\hat{\mathcal{M}}$ is expressed as follows.
^$\vspace{-0.2cm}\boxed{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\hat{\mathcal{M}}}=\mathcal{M}-\lambda*{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{H}_{v}}}$ (7)
where $\lambda$ is a hyperparameter.
4 Language and dataset
Figure 3: Results on the MultiJail dataset. Red bars represent the base model’s unsafe outputs, while blue bars denote outputs from the safe model Soteria. Languages are categorized by resource availability: H (high resource), M (mid resource), and L (low resource). The substantial reduction in unsafe content across high-, mid-, and low-resource languages highlights the effectiveness of the Soteria compared to the base model. The ASR values presented here range from 0 to 1. To express them as percentages, simply multiply by 100. Lower is better.
Languages: Following Deng et al. (2024a), we consider twelve languages across high-, medium- and low-resource categories. From the high-resource language category, we consider English (En), Chinese (Zh), German (De), French (Fr), and Spanish (Es). For the medium-resource language category, Arabic (Ar), Thai (Th), Bulgarian (Bg), and Hindi (Hi). For low-resource language category, we include Tamil (ta), Bengali (bn), and Telugu (te).
Datasets: We assess Soteria using two established datasets, MultiJail Deng et al. (2024b) and XSafety Wang et al. (2024). In addition, we introduce a new multilingual safety dataset XThreatBench, constructed based on the policy violations outlined by Meta Qi et al. (2023a). A detailed description of each dataset follows. We include the dataset details of XSafety and the corresponding experimental results in the Appendix A due to space constraints.
MultiJail: This dataset is the first multilingual translated jailbreak benchmark designed to assess the safety vulnerabilities of large language models across multiple languages. It contains 3150 manually translated queries across 10 languages, covering high-resource (English, Chinese, Italian, Vietnamese), medium-resource (Arabic, Korean, Thai), and low-resource (Bengali, Swahili, Javanese) languages. Built from harmful queries in the GPT-4 report OpenAI et al. (2024) and Anthropic’s red-teaming dataset Ganguli et al. (2022), it explores unintentional and intentional jailbreaks, where translation itself serves as a jailbreak method. For our experiments, we use google translate [^2] https://translate.google.com to translate English queries into other languages when they are not present in the dataset.
XThreatBench: We propose a multilingual safety benchmark on general harm designed to rigorously evaluate LLM vulnerabilities across 10 high-risk categories, including sexual content, child exploitation, economic fraud, hate speech, illegal activities, cyber threats, physical harm, political manipulation, privacy violations, and deception. XThreatBench features 3,000 adversarial prompts across 12 languages ensuring native linguistic authenticity with generic harm. Each category contains 25 instances, systematically crafted to test LLM safety mechanisms against general attack scenarios, evasive jailbreak tactics, and multilingual exploits. To maximize adversarial quality, we implemented a three-stage verification process, combining partial human intervention, GPT-4 based filtering, and Perspective API based toxicity scoring (retaining only queries with a toxicity score of 0.75+). Unlike existing benchmarks, XThreatBench adheres to safety policies while offering a robust, multilingual evaluation framework for assessing LLM safety in high-risk, adversarial environments.
5 Experimental setup
In this section, we first introduce the language models used in our evaluation, selected for their multilingual capabilities and diverse linguistic distributions. Next, we define our evaluation metric, attack success rate (ASR), to quantify safety violations. Subsequently, we describe the jailbreak attack baselines. To benchmark our proposed safety mechanism, we compare it against existing English language-centric safety alignment approaches.
5.1 Language models
Figure 4: Results on the XThreatBench dataset. Red bars represent the base model’s unsafe outputs, while blue bars denote outputs from the safe model Soteria. Languages are categorized by resource availability: H (high-resource), M (mid-resource), and L (low-resource). The substantial reduction in unsafe content across high-, mid-, and low-resource languages highlights the effectiveness of Soteria compared to the base model. The ASR values presented here range from 0 to 1. To express them as percentages, simply multiply by 100. Lower is better.
We use four open-weight language models to evaluate our proposed safety mechanism. We consider instruction-tuned versions of llama3.1-8b-instruct Grattafioriet et al. (2024), Qwen2-7B-Instruct Yang et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Phi-3.5-mini-instruct Abdin et al. (2024). We select these models because they are widely used instruction-tuned LLMs with diverse multilingual capabilities. They perform well in high-resource languages such as English and Chinese, while their support for medium- and low-resource languages varies. This selection allows us to evaluate safety mechanisms across different linguistic distributions and understand their effectiveness in multilingual settings.
5.2 Language specific tasks used
Following the tasks outlined in Todd et al. (2024), we use five datasets: (1) agnews, a text classification dataset of news headlines and early article sentences labeled as business, science/technology, sports, or world; (2) antonym, where each input word requires generating its opposite; (3) country-capital, comprising 197 country-capital pairs and requiring the capital city given a country; (4) country-currency, listing countries alongside their currencies; and (5) sentiment, which contains 1167 movie review sentences labeled as either positive or negative. Further, we translate these task datasets into specific languages and identify the attention heads.
5.3 Evaluation metrics
Following previous studies Chao et al. (2024); Qi et al. (2023b), we assess the safety standards of language models using the attack success rate (ASR). ASR measures the proportion of harmful responses generated by the model out of the total number of harmful queries and is defined as follows.
ASR=# responses not aligned with Safety# input queries to $\text{ASR}=\frac{\text{\# responses not aligned with {Safety}}}{\text{\# inputqueries to LLM}}$
We classify a response as harmful or unethical if it violates the usage policies of OpenAI [^3] https://openai.com/policies/usage-policies and Meta [^4] https://ai.meta.com/llama/use-policy.
5.3.1 GPT-4o as a safety evaluator
To reliably and efficiently assess a model’s safety performance, we employ GPT-4o as the annotator, leveraging its strong alignment with human judgments on harmful responses Qi et al. (2023b); Banerjee et al. (2024b); Hazra et al. (2024a). In particular, we adopt the evaluation prompt [^5] Given the following text, analyze and assess its ethical implications. Determine whether the content adheres to ethical standards or if there are elements that raise ethical concerns. proposed by Banerjee et al. (2024b). Cases in which GPT-4o declines to provide annotations due to content filtering are excluded from our calculations. A lower ASR score indicates a safer model.
5.4 Baselines
We compare our multilingual safety mechanism with the following safety mechanism techniques, which involve language model parameters. The descriptions of these baselines are as follows.
(1) Safety arithmetic Hazra et al. (2024a): The safety arithmetic framework improves model safety across the base, fine-tuned, and edited models where safety risks emerge due to fine-tuning artefacts, or unintended editing consequences. It adjusts model parameters and realigns the latent space to reduce harmful outputs and ensure safer content generation.
(2) Resta Bhardwaj et al. (2024): This method restores safety in fine-tuned LLMs by adding a safety vector from the difference between a safety-aligned and an unaligned model. It counteracts safety degradation from supervised fine-tuning and enhances alignment using drop and rescale (DARE) Yu et al. (2024) to remove redundant delta parameters before applying Resta.
(3) TIES Yadav et al. (2023): In this method, we consider the top 3% of parameters in the harm vector $H_{v}$ and then subtract the trimmed harm vector from the target language model.
(3) Self-defense Deng et al. (2024b): We could not compare the self-defense method which suggests that simple fine-tuning with a specific dataset can restore multilingual safety, due to the unavailability of the dataset mentioned in the paper.
Figure 5: Comparison of Soteria with other baselines [^7] We define average of High resources as High, and similarly for Mid and Low. This also holds for Figure 6 and Table 2..
En Zh Es Fr De Hi Ar Th Bg Bn Ta Te
High resource Mid resource Low resource
Lang B SU B SU B SU B SU B SU B SU B SU B S B SU B SU B SU B SU
Multijail
Llama 3.1 0.43 0.26 0.51 0.2 0.37 0.2 0.41 0.1 0.36 0.19 0.54 0.22 0.32 0.23 0.49 0.34 0.39 0.2 0.34 0.32 0.52 0.22 0.3 0.16
Qwen 2 0.35 0.25 0.23 0.1 0.13 0.11 0.2 0.04 0.23 0.06 0.37 0.2 0.08 0.08 0.26 0.08 0.15 0.1 0.14 0.11 0.47 0.34 0.3 0.28
Mistral v3 0.35 0.12 0.37 0.08 0.2 0.19 0.27 0.19 0.29 0.22 0.27 0.18 0.32 0.28 0.33 0.28 0.25 0.17 0.2 0.02 0.1 0.04 0.05 0.02
Phi 3.5 0.21 0.04 0.22 0.04 0.18 0.1 0.25 0 0.16 0.04 0.35 0.2 0.21 0.18 0.21 0.2 0.19 0.14 0.16 0.15 0.26 0.22 0.23 0.21
XThreatBench
Llama 3.1 0.21 0.13 0.25 0.18 0.22 0.12 0.18 0.1 0.21 0.1 0.17 0.17 0.29 0.23 0.23 0.13 0.29 0.22 0.28 0.18 0.2 0.19 0.13 0.11
Qwen 2 0.14 0.09 0.12 0.04 0.12 0.09 0.11 0.05 0.1 0.06 0.14 0.13 0.15 0.1 0.18 0.18 0.14 0.1 0.13 0.13 0.22 0.22 0.18 0.13
Mistral v3 0.16 0.1 0.26 0.13 0.18 0.04 0.23 0.18 0.16 0.16 0.26 0.15 0.3 0.26 0.24 0.23 0.3 0.14 0.25 0.08 0.06 0.02 0.05 0
Phi 3.5 0.07 0.02 0.12 0.12 0.09 0.07 0.14 0.07 0.06 0.05 0.13 0.11 0.14 0.18 0.05 0.16 0.14 0.16 0.14 0.17 0.1 0.06 0.12 0.18
Table 1: Results from SoteriaU. We identify functional neurons by selecting the majority of heads across all languages and then retaining 50% of the most significant heads. B: base model, SU: SoteriaU. Green = lower, blue = equal, red = higher vs. base model.
6 Main results
Here we demonstrate the results from Soteria across different languages in Figure 3 and Figure 4.
Results for different datasets:
MultiJail: Evaluation of our proposed method Soteria across multiple language models demonstrates substantial disparities in adversarial robustness across high-resource, medium-resource, and low-resource languages (see Figure 3). For high-resource languages, the ASR is moderately high, with Llama 3.1 and Qwen 2 exceeding 50% ASR in certain languages. However, after applying Soteria, ASR is reduced by 40–60%, with En and Es showing the most substantial reductions, dropping to nearly 20–25% ASR in the safe models. Zh, however, exhibits a less consistent decline, with some models retaining ASR levels above 30%, indicating that adversarial robustness is still incomplete for logographic scripts. For medium-resource languages , ASR reductions are less pronounced compared to high-resource languages. The base model’s ASR for these languages is often higher than 50%. After applying our safety mechanisms, the ASR drops by approximately 30–50%, with the most effective reductions observed in Hn and Bg, where ASR reaches 25–35% post-safety alignment. Notably, Mistral 0.3 and Phi 3.5 outperform Llama 3.1 and Qwen 2 in these languages, with ASR reductions exceeding 50% in some cases.Low-resource languages present the greatest challenge, as their baseline ASR is the highest among all language groups, often exceeding 60%. Despite safety interventions, ASR reductions are minimal, typically ranging between 15–30%. Even in the best-performing models, the final ASR rarely drops below 40%. Llama 3.1 and Qwen 2 struggle the most, with ASR remaining as high as 50% even after applying our safety mechanism. In contrast, Mistral 0.3 and Phi 3.5 achieve slightly better reductions but still maintain ASR levels around 35–45%.
XThreatBench: In case of this dataset (see Figure 4), the evaluation of ASR across different language models reveals notable variations in vulnerability before and after the application of Soteria. In high-resource languages, base models exhibit ASR values ranging from approximately 25–35%, with Llama 3.1 and Qwen 2 showing the highest susceptibility. Post-safety interventions, ASR is reduced significantly to 5–15%, demonstrating the efficacy of the mitigation strategies. In medium-resource languages, initial ASR ranges between 20–40%, with Mistral 0.3 showing comparatively lower vulnerability. After applying Soteria, ASR declines to 10–20%, though the reduction is less pronounced than in high-resource languages. Low-resource languages remain the most vulnerable, with base ASR values between 25–30%, and post-safety using Soteria, ASR still hovering around 10–20%, indicating persistent risks despite intervention. Among all models, Phi 3.5 consistently demonstrates the lowest post-safety ASR across all language groups, staying within 5%–15%.
General capabilities: We evaluate our framework’s impact on overall model capabilities using utility tests (MMLU Hendrycks et al. (2021) 5-shot and TruthfulQA Lin et al. (2022)). The results closely mirror each base model’s performance. For the safe version of Llama 3.1, we observe the MMLU performance at 72.9 (vs. 73 from the baseline), and TruthfulQA at 44.14 (vs. 44.14 for the baseline). The safe version of Qwen exactly matched its base values (70.3, 54.2). Mistral yielded 61.79 MMLU (vs. 61.84) and 59.34 TruthfulQA (vs. 59.37), while Phi also retained its baseline scores of 69 (MMLU) and 64 (TruthfulQA).
Comparison with the baselines: We compare Soteria with three English-centric safety alignment methods as discussed above – safety-arithmetic, Resta, and TIES – by examining the ASR values for high-, medium-, and low-resource languages. Figure 7 presents the results for two models, Llama 3.1 and Qwen 2, using the Multijail and XThreatBench datasets. Across all baselines, Soteria consistently achieves the lowest ASR. On Llama 3.1 with the Multijail dataset, the baseline method’s ASR ranges from 30–40% in high-resource languages, while for Soteria it is about 15–20%. Both TIES and Resta provide moderate decreases (30–35%), and safety-arithmetic does slightly better (25–30%). However, Soteria consistently outperforms these methods by 5–10%. Similar trends hold for medium- and low-resource languages. A comparable trend is also observed from Qwen 2. For Multijail, the baseline ASR is approximately 28–30% in high-resource settings, whereas TIES, Resta, and safety-arithmetic reduce it to 20–25%. Soteria pushes the ASR even lower, to around 15–20%. These findings also generalize to XThreatBench, reinforcing the robustness of Soteria across diversely resourced languages, models and datasets.
7 Language universals
[…]
Results: We observe that the SoteriaU consistently produces lower ASR compared to three base models across all tested languages and model backbones (see Table 1). For example, for the Multijail dataset, Llama 3.1’s ASR in English drops from 43% (base) to 26% (safe), while in Chinese it decreases from 51% to 20%. Similar reductions are observed for Qwen 2 (35% to 25% in English), Mistral 0.3 (35% to 12% in English), and Phi 3.5 (21% to 4% in English), demonstrating that SoteriaU effectively curtails harmful responses. This pattern persists for the XThreatBench dataset as well, where the safe configurations again achieve notably lower ASRs across languages (e.g., Phi 3.5’s English ASR goes from 7% to 2%). In the mid-resource languages like Arabic in Multijail, Llama 3.1’s ASR drops from 32% to 23%, while in low-resource Tamil, it decreases from 52% to 22%. Across both the Multijail and XThreatBench datasets, SoteriaU consistently outperforms the base models by lowering harmful outputs in a language-agnostic manner. These results highlight the robustness and effectiveness SoteriaU, regardless of whether the language is high-, mid or low-resourced.
Figure 6: Trade-off between ASR and % heads probed.
8 LLM jailbreaks
We employ recent jailbreak methods to evaluate the robustness of Soteria.
POATE Sachdeva et al. (2025): The POATE jailbreak method manipulates LLMs using contrastive reasoning, subtly reframing harmful queries into their opposites. Unlike direct exploits, it combines adversarial templates to bypass safety measures and trigger unintended responses.
Refusal direction Arditi et al. (2024): LLMs’ refusal behaviour follows a single identifiable direction in activation space. Removing this refusal direction (RDR) bypasses safety measures, enabling harmful responses, while adding it increases refusals. This discovery led to a white-box jailbreak method using a rank-one weight modification to disable refusals with minimal impact on other functions.
Results: For both the MultiJail and XThreatBench evaluations for the Llama 3.1 8B model, our strategy consistently yields lower ASR than the baseline jailbreaks, indicating a substantial reduction in the model’s vulnerability (see Table 2). In MultiJail, POATE’s high threat setting decreases from 0.53 to 0.33, and RDR drops from 0.49 to 0.29. Mid and low threat scenarios show similar improvements. In XThreatBench, the reduction is even more pronounced: POATE’s high threat rate falls from 0.46 to 0.13 and RDR goes from 0.30 to 0.11. These results demonstrate that Soteria significantly mitigates the impact of advanced jailbreak techniques across all threat levels for Llama 3.1 8B [^8] Results are similar for other models and are not shown due to paucity of space..
9 ASR vs. % heads probed
Figure 6 shows how the ASR changes as we vary the percentage of attention heads in the model, for three different resource settings. All three settings initially exhibit their highest ASRs at 25% heads, suggesting that using only a small fraction of heads leaves the model more vulnerable. When the percentage of heads increases to 50%, ASRs drop noticeably across the board, indicating a clear gain in robustness at this midpoint. If we use more than 50% heads, increasingly smaller improvement rates are observed. This shows that after a certain point, adding more heads brings less benefit. Assuming that each layer in a 8B model has $\sim 32$ heads and there are $\sim 32$ such layers, we need to probe $0.5\times 32\times 32=512$ heads. Further the dimension of the corresponding projection matrix $W^{O}_{li}$ is $\sim 4096\times 128$. Thus, roughly the % of heads probed is only $\left(\frac{512(heads)\times 128(dimension)\times 4096(params)}{8\mathrm{B}}\right)\times 100\sim 3\%$
High Mid Low
MultiJail
Base-J S-J Base-J S-J Base-J S-J
POATE 0.53 0.33 0.61 0.36 0.62 0.36
RDR 0.49 0.29 0.53 0.30 0.61 0.36
XThreatBench
POATE 0.46 0.13 0.45 0.18 0.44 0.19
RDR 0.30 0.11 0.39 0.16 0.37 0.16
Table 2: Robustness of Soteria against SOTA jailbreak attacks. S-J: Soteria.
10 Conclusion
We introduce Soteria, a lightweight yet powerful safety alignment method that fine-tunes language-specific “functional neurons” in multilingual LLMs. By adjusting only a fraction of parameters, Soteria effectively curbs policy violations across high-, mid-, and low-resource languages without compromising overall performance. Our XThreatBench dataset, derived from real-world policy violations, demonstrates that this targeted parameter steering outperforms baseline safety approaches. These results highlight the value of language-aware interpretability and the practicality of scalable multilingual safeguards, advancing inclusive and ethically responsible AI.
11 Limitation
A key limitation of Soteria lies in its reliance on per-language functional neuron identification, which requires accurate language segmentation and task-based data in each target language. In practice, resource constraints, limited training data, and complexities in script variation or morphology can reduce the precision of head selection. Moreover, although Soteria improves safety across many languages, it does not guarantee comprehensive coverage of every cultural nuance or emergent harmful behavior.
12 Ethical Consideration
In designing and evaluating Soteria, we prioritized responsible data use and clear ethical practices: XThreatBench was curated exclusively from synthetic or publicly available prompts crafted to evaluate harmful scenarios without including any personal or sensitive user data. We aligned our methodology with widely recognized industry norms, ensuring minimal data collection and protecting user privacy. Moreover, we respected the cultural nuances that shape perceptions of harm by incorporating broad content moderation principles from organizations like Meta and OpenAI. By balancing robust multilingual safety mechanisms with careful attention to legitimate expression and cultural diversity, our approach aims to foster a more secure yet equitable AI environment.
Appendix A Additional experiment
High Resource Mid Resource Low Resource
En Zh De Fr Es Bg Hi Th Ar Bn Te Ta
Languages B S B S B S B S B S B S B S B S B S B S B S B S
llama3.1-8b-instruct 0.12 0.05 0.14 0.07 0.12 0.03 0.09 0.03 0.08 0.01 0.17 0.08 0.12 0.05 0.11 0.05 0.09 0.06 0.13 0.08 0.11 0.07 0.13 0.08
Qwen2-7B-Instruct 0.08 0.05 0.03 0.02 0.04 0.03 0.04 0.02 0.03 0.02 0.05 0.02 0.06 0.05 0.04 0.03 0.03 0.02 0.07 0.04 0.07 0.07 0.09 0.08
Mistral-7B-Instruct-v0.3 0.11 0.03 0.1 0.02 0.08 0.04 0.1 0.06 0.06 0.03 0.09 0.05 0.11 0.05 0.08 0.06 0.08 0.1 0.08 0.02 0.04 0.01 0.02 0.01
Phi-3.5-mini-instruct 0.08 0.01 0.11 0.05 0.06 0.02 0.09 0.03 0.06 0.02 0.07 0.06 0.09 0.05 0.08 0.06 0.09 0.07 0.04 0.03 0.05 0.05 0.02 0.02
Table 3: Results on the XSafety dataset. B represent the base model’s unsafe outputs, while S denote outputs from Soteria. The substantial reduction in unsafe content across high-, mid-, and low-resource languages highlight the effectiveness of the Soteria compared to the base model. Lower is better. Green = lower, blue = equal, red = higher vs. base model.
High Resource Mid Resource Low Resource
En Zh De Fr Es Bg Hi Th Ar Bn Te Ta
Languages B S B S B S B S B S B S B S B S B S B S B S B S
llama3.1-8b-instruct 0.12 0.06 0.14 0.11 0.12 0.07 0.09 0.04 0.08 0.03 0.17 0.09 0.12 0.07 0.11 0.07 0.09 0.04 0.13 0.12 0.11 0.05 0.13 0.08
Qwen2-7B-Instruct 0.08 0.06 0.03 0.03 0.04 0.01 0.04 0.02 0.03 0.03 0.05 0.03 0.06 0.04 0.04 0.02 0.03 0.03 0.07 0.05 0.07 0.04 0.09 0.04
Mistral-7B-Instruct-v0.3 0.11 0.02 0.1 0.1 0.08 0.01 0.1 0.04 0.06 0.05 0.09 0.09 0.11 0.06 0.08 0.1 0.08 0.1 0.08 0.02 0.04 0 0.02 0.01
Phi-3.5-mini-instruct 0.08 0.01 0.11 0.04 0.06 0.03 0.09 0.01 0.06 0.04 0.07 0.06 0.09 0.07 0.08 0.09 0.09 0.09 0.04 0.04 0.05 0.04 0.02 0.02
Table 4: Results from Soteria. We identify functional neurons by selecting the majority of heads across all languages and then retaining 50% of the most significant heads. B: base model, S: Soteria. Green = lower, blue = equal, red = higher vs. base model.
XSafety: This is a multilingual safety benchmark designed to evaluate LLMs across multiple languages. It consists of 2,800 manually translated instances covering 14 safety categories in 10 widely spoken languages: English, Chinese, Spanish, French, Bengali, Arabic, Hindi, Russian, Japanese, and German. Built from existing monolingual safety datasets, XSafety was translated and verified by annotators, ensuring cross-lingual consistency. The benchmark reveals significant safety gaps in non-English responses, emphasizing the need for multilingual safety alignment. For our experiments, we use google translate [^9] https://translate.google.com to translate English queries into other languages when they are not present in the dataset.
A.1 Result for XSafety dataset
Figure 7: Identified top 20 heads for Llama 3.1 8B for all languages.
The results presented in Table 3 illustrate the substantial improvements achieved by integrating the Soteria framework across a wide range of languages and language models. The comparison between the baseline models (B) and the safe models (S) reveals a significant reduction in unsafe outputs across high-, mid-, and low-resource languages. This consistent improvement underscores the effectiveness of Soteria as a robust and scalable solution for mitigating unsafe content generation in multilingual LLMs.
In high-resource languages such as English, Chinese, German, French, and Spanish, the impact of Soteria is particularly noteworthy. For example, in English, the unsafe output rate for the Llama 3.1 model drops from 0.12 in the baseline to 0.05 with Soteria. Similar improvements are observed in Chinese (0.14 to 0.07) and German (0.12 to 0.03), reflecting a substantial reduction in unsafe behavior. The safe versions of models like Qwen 2 and Mistral show comparable improvements, with Qwen 2 reducing the unsafe rate in Chinese from 0.03 to 0.02 and Mistral achieving a reduction in English from 0.11 to 0.03. These results demonstrate that Soteria not only improves safety for individual models but also generalizes effectively across different architectures and languages.
Mid-resource languages such as Bulgarian, Hindi, Thai, and Arabic pose additional challenges due to their relatively limited training data. Despite these difficulties, Soteria delivers significant reductions in unsafe outputs across all models. For instance, in Bulgarian, the unsafe rate for Llama 3.1 drops from 0.17 to 0.08, a nearly 50% improvement. Similar trends are seen in Hindi, where the rate falls from 0.12 to 0.05, and Thai, with a reduction from 0.11 to 0.05. Qwen 2 also demonstrates strong performance improvements in these languages, particularly in Hindi, where it reduces the unsafe rate to 0.05. Even in Arabic, which presents unique challenges, models like Mistral and Phi 3.5 achieve remarkably low unsafe rates, indicating that Soteria is effective in maintaining safety across diverse linguistic and cultural contexts.
The performance of Soteria in low-resource languages such as Bengali, Telugu, and Tamil further validates its adaptability and scalability. Low-resource languages often exhibit higher baseline unsafe output rates due to their underrepresentation in training data. However, Soteria consistently reduces these rates, demonstrating its capacity to address safety concerns in less-resourced linguistic settings. In Bengali, for example, Llama 3.1 reduces the unsafe rate from 0.13 to 0.08, while Telugu and Tamil see similar improvements, with reductions from 0.11 to 0.07 and 0.13 to 0.08, respectively. Notably, Mistral and Phi 3.5 continue to perform exceptionally well, with Mistral achieving an impressively low unsafe rate of 0.01 in Tamil.
The results presented across these language groups make it clear that Soteria offers a transformative approach to improving safety in large language models. The consistent reductions in unsafe outputs, ranging from high-resource to low-resource languages, highlight the robustness and generalizability of the framework.
A.2 XSafety (language universal)
In Table 4 for high-resource languages such as English, Chinese, German, French, and Spanish, the reduction in unsafe outputs is substantial. For example, in English, the unsafe rate for Llama 3.1 drops from 0.12 to 0.06, and in German, it declines from 0.12 to 0.07. Similar improvements are observed across other high-resource languages. Qwen 2 reduces the unsafe rate in French from 0.04 to 0.02 and shows consistent gains across other languages like Chinese and Spanish. Mistral stands out in English, where it brings down the unsafe rate from 0.11 to 0.02. These reductions reflect the precision with which Soteria identifies and mitigates unsafe content while maintaining the language models’ core functionality.
The mid-resource languages – Bulgarian, Hindi, Thai, and Arabic – further illustrate Soteria’s adaptability. Bulgarian, for instance, sees a significant improvement with Llama 3.1 reducing the unsafe rate from 0.17 to 0.09, and Hindi experiences a similar reduction from 0.12 to 0.07. Mistral also achieved substantial progress in Bulgarian, reducing unsafe outputs to 0.09. These results are a clear indicator that Soteria effectively addresses the unique challenges presented by languages with moderately available resources, ensuring more controlled output across different linguistic patterns and complexities.
In low-resource languages such as Bengali, Telugu, and Tamil, where limited data often results in higher baseline unsafe rates, Soteria continues to deliver meaningful reductions. Llama 3.1 reduces the unsafe rate in Bengali from 0.13 to 0.08, while Telugu sees an improvement from 0.11 to 0.05. Tamil shows equally promising results, with multiple models significantly lowering unsafe outputs. Notably, Mistral reduces the unsafe rate in Tamil to 0.01, demonstrating that Soteria can extend its impact even to data-scarce settings without requiring extensive retraining or language-specific adjustments.
Overall, the results highlight Soteria’s capacity to improve model safety at scale, offering a practical and efficient approach to reducing unsafe outputs across languages with diverse resource levels. The consistent reduction in unsafe rates across models and languages indicates that Soteria is not only scalable but also robust in its generalization across linguistic and cultural boundaries.
Appendix B Attention head patterns and their implications
One intriguing characteristic of LLMs is how their top-valued language‐specific attention heads tend to cluster by resource level of the language. Analyses of a smaller-parameter model (e.g., Llama 3.1 8B‐parameter variant) reveal that high‐resource languages (such as English, Chinese, Spanish, German, and French) and mid‐resource languages (such as Hindi, Arabic, Thai, and Bulgarian) exhibit peak attention heads in roughly the same mid‐level layers (e.g., layers 12–20 with head indices 16–24). Meanwhile, for low‐resource languages the strongest attention heads manifest in later layers (e.g., layers 28–31 with head indices 15–23) (see Figure 7).
(1) Language-specific universal heads: Despite the differences in where each language’s top heads appear, some heads consistently contribute to cross‐lingual understanding – the so‐called “universal” heads. Identifying and enhancing these universal heads can make the model’s latent space more cohesive across languages, improving zero‐shot or few‐shot performance for underrepresented languages.
(2) Future directions: Beyond raw performance, attention‐head analysis also provides new insights to tackle task-specific attention heads, misalignment, and hallucination issues. If certain heads consistently carry problematic correlations, shifting or refining their latent space (“steer them to a safe side”) can enhance overall alignment and trustworthiness.
These findings underscore the delicate interplay between multilingualism and architectural depth in multilingual models. By homing in on the most influential heads and understanding why they appear where they do, we gain powerful levers for improving cross‐lingual performance, minimizing unsafe content generation, and facilitating more robust language support, even for the world’s most resource sparse tongues.